2x +3y=10
6x+4y=5
Adiition (elimination method)

Answer by checkley77(12569)

2x +3y=10 MULTIPLY BY -3 & ADD.
6x+4y=5
-6X-9Y=-30
------------------
-5Y=-25
Y=-25/-5
Y=5 ANS.
2X+3*5=10
2X+15=10
2X=10-15
2X=-5
X=-5/2
X=-2.5 ANS.
PROOF:
6*-2.5+45=50
-15+45=30
30=30